OpenStudy (yolo_king):

I need help with a Riddle There 12 pens on a table. You take 5, return 3, and you take 2. How many do you have left?

OpenStudy (isaiah.feynman):

2

OpenStudy (dumbcow):

this isnt a riddle haha 12 -5 = 7 7+3 = 10 10-2 = 8

OpenStudy (dqswag):

8

OpenStudy (yolo_king):

I submitted 8 and i got it wrong :/

OpenStudy (isaiah.feynman):

@dumbcow Not in that manner.

OpenStudy (dqswag):

12-5+3-2=8

OpenStudy (yolo_king):

I think 2 is about right

OpenStudy (dqswag):

Lol this aint no redile

OpenStudy (dumbcow):

oh sorry there are 8 left on table .... thus you have 4

OpenStudy (isaiah.feynman):

Its not how many left on the table, its how many you have with you.

OpenStudy (yolo_king):

Ahh i see

OpenStudy (dqswag):

ohhhhhhhh

OpenStudy (dqswag):

lol ima noob

OpenStudy (yolo_king):

4 Is rigth :D Thx

OpenStudy (isaiah.feynman):

Oh I mean 4.